# Quantum Incompatibility in Parallel versus Antiparallel Spins

### v2

@@ I Proof of Proposition 1
So far we have analyzed joint measurability for unsharp spin measurements only. However, the notion can be generalized for more general POVMs. A ‘$K$’ outcome qubit POVM $\mathcal{A}$ is the collection of $K\penalty 10000\ (<\infty)$ effects adding up to identity operator. Formally
@@ II Enhanced measurement compatibility in GPT
With this we are now in a position to state our next result.
Theorem 2S.
For any set of qubit POVM measurements $\{\mathcal{A}_{l}\}_{l=1}^{N}$, their joint measurability on the $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory implies their joint measurability on the $\uparrow\hskip-1.42271pt\uparrow$ configuration in minimal tensor product GPT.
@@
In Theorem 2 of our main manuscript we have shown that the observables $\mathrm{SIC}_{4}$ are compatible in quantum theory on $\uparrow\hskip-1.42271pt\downarrow$ configuration. Now as a corollary of Theorem 2S it follows that these observables are compatible on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT.
@@
One might expect that the converse of Theorem 2S also holds true, namely joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT would imply their joint measurability on $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory. However, as we argue this is not the case in general. Joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT implies existence of a collection of operators $\left\{\Pi_{\bf r}\in\mathbb{E}^{AB}_{\min}\penalty 10000\ |\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\right\}_{\bf r}$, where some of the $\Pi_{\bf r}$’s could be entangled. Consequently, the respective $\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)$’s are not positive operators, and hence do not constitute a valid measurement in quantum theory.

### v3

@@ I Proof of Proposition 1
The notion joint measurability is in general studied for general POVMs. A ‘$K$’ outcome qubit POVM $\mathcal{A}$ is the collection of $K\penalty 10000\ (<\infty)$ effects adding up to identity operator. Formally
@@ II Enhanced measurement compatibility in GPT
With this we are now in a position to prove the Theorem 4 of main manuscript.
@@
In Theorem $2$ of our main manuscript we have shown that the observables $\mathrm{SIC}_{4}$ are compatible in quantum theory on $\uparrow\hskip-1.42271pt\downarrow$ configuration. Now as a corollary of Theorem 4 it follows that these observables are compatible on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT.
@@
One might expect that the converse of Theorem 4 also holds true, namely joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT would imply their joint measurability on $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory. However, as we argue this is not the case in general. Joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT implies existence of a collection of operators $\left\{\Pi_{\bf r}\in\mathbb{E}^{AB}_{\min}\penalty 10000\ |\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\right\}_{\bf r}$, where some of the $\Pi_{\bf r}$’s could be entangled. Consequently, the respective $\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)$’s are not positive operators, and hence do not constitute a valid measurement in quantum theory.